OpenStudy (anonymous):

Mary invested $12000.00 and got 1% profit in fund a! fund b she made 6% her total profit was 3% how much did she invest in fund b

OpenStudy (tkhunny):

It's just a translation problem. a + b = 12000 a*0.01 + b*0.06 = 12000*0.03

OpenStudy (calculusxy):

Which you can therefore make it to0.01a+0.06b=360

OpenStudy (mathstudent55):

From the wording of the problem, I don't understand if the $12,000 was the investment only in fund A or if the $12,000 was the total investment in funds A and B.

OpenStudy (anonymous):

The problem said the investment for fund a was 12000.00 it did not give an amount for fund b. just said she earned 6% profit And the total profit for both funds was 3%

OpenStudy (tkhunny):

Wow! 12000*0.01 b*0.06 (12000+b)*0.03 Go!

OpenStudy (anonymous):

Finally asked for the solution and they said $6,000 was the amount of fund b. have no idea how to get that answer

OpenStudy (mathstudent55):

Since the problem can be interpreted as $12,000 being only for Fund A or for the total of both funds, I solved it both ways below. Let's call the amount invested in Fund A, a, and the amount invested in Fund B, b. 1. If the total amount invested is $12,000 a + b = 12000 a = 12000 - b Interest from Fund A Interest from Fund B Total Interest 0.01(12000 - b) + 0.06b = 0.03(12000) When you solve for b, you get b = 4800 That means the money invested in Fund B is $4800 2. If the entire $12,000 was invested in Fund A, and additional money was invested in Fund B Interest from Fund A Interest from Fund B Total Interest 0.01(12000) + 0.06b = 0.03(12000 + b) When you solve for b, you get b = 8000. That means the money invested in Fund B is $8000 I don't see how you get $6000 invested in Fund B.